### v1

# Self-Organising Memristive Networks as Physical Learning Systems

Abstract
Learning with physical systems is an emerging paradigm that seeks to harness the intrinsic nonlinear dynamics of physical substrates for learning. The impetus for a paradigm shift in how hardware is used for computational intelligence stems largely from the unsustainability of artificial neural network software implemented on conventional transistor-based hardware. This Perspective highlights one promising approach using physical networks comprised of resistive memory nanoscale components with dynamically reconfigurable, self-organising electrical circuitry. Experimental advances have revealed the non-trivial interactions within these Self-Organising Memristive Networks (SOMNs), offering insights into their collective nonlinear and adaptive dynamics, and how these properties can be harnessed for learning using different hardware implementations. Theoretical approaches, including mean-field theory, graph theory, and concepts from disordered systems, reveal deeper insights into the dynamics of SOMNs, especially during transitions between different conductance states where criticality and other dynamical phase transitions emerge in both experiments and models. Furthermore, parallels between adaptive dynamics in SOMNs and plasticity in biological neuronal networks suggest the potential for realising energy-efficient, brain-like continual learning. SOMNs thus offer a promising route toward embedded edge intelligence, unlocking real-time decision-making for autonomous systems, dynamic sensing, and personalised healthcare, by enabling embedded learning in resource-constrained environments. The overarching aim of this Perspective is to show how the convergence of nanotechnology, statistical physics, complex systems, and self-organising principles offers a unique opportunity to advance a new generation of physical intelligence technologies.
1 Introduction
Physical learning systems can be broadly described as physical substrates whose intrinsic nonlinear dynamics can be exploited for learning [1, 2]. Examples include analog electrical circuits, photonic and opto–electronic systems, as well as stochastic spintronic systems [3]. A particularly promising approach aims to harness native brain–like, or neuromorphic, properties exhibited by certain nanoscale electronic devices, such as memristors (see Box 1), with physical memory and neuromorphic dynamics [4]. Among the most promising examples of such systems are Self-Organising Memristive Networks (SOMNs). These physical networks are described as self–organising because under external electrical stimulation, they can spontaneously acquire specific novel structural and functional properties [5], effectively making them dynamically reconfigurable electrical circuits. These macroscale, observable properties emerge from the collective dynamical interactions between their constituent nanoscale elements, in the spirit of “More is Different” [6]. As highlighted in this Perspective, SOMNs are unique physical learning substrates because their ability to encode, store, and process information arises directly from their material and structural properties, positioning them at the intersection of condensed matter physics, neuromorphic computing, nanotechnology, and network science.
The importance of the structure and connectivity of biological neuronal networks to learning was recognised by some of the earliest work on artificial intelligence [7, 8, 9]. Rosenblatt’s perceptron introduced probabilistic learning and statistical separability as central to cognition; he envisioned the perceptron as a biologically–inspired model for emergent dynamics from random but structured physical wiring, while also noting practical limitations in scaling such architectures [8, 9, 10, 11]. These early insights motivated the first experimental realisations of SOMNs as scalable physical networks with neuromorphic architectures and dynamics achieved using disordered nano-electronic materials [12, 13, 14, 15, 16].
An active area of research in neuromorphic computing is the exploration of physical substrates that can be engineered to mimic the dynamics and efficiency of biological neurons and synapses [17]. In particular, memristor devices connected in a cross-bar array architecture map directly onto the bipartite graph structure of artificial neural networks (ANNs) and can thus serve as a physical instantiation of matrix–vector multiplications for efficient inference of ANN models [18, 19].
[…]
In this Perspective, we posit that SOMNs are unique physical learning substrates because they mimic the structure–function relation of the brain, which is the archetypal non–equilibrium learning system [22, 23, 24, 25, 26, 27, 28]. Typically comprised of nanowires [12, 13, 20] or nanoparticles [29, 15, 16], SOMNs are formed via bottom–up synthesis, which naturally produces a high density of interconnections, far surpassing the limits of conventional top-down lithographic fabrication [14]. Their electronic architecture is also inherently stochastic [30], just like biological neuronal networks, i.e. no two single brains are identical.
Figure 1 summarises the key features of SOMNs as physical learning systems. In contrast to conventional hardware computing systems (and many unconventional ones), SOMNs can process information efficiently without the need for direct addressability of individual components [1], similarly to biological neural networks. SOMNs can operate as physical learning systems by applying relatively simple learning schemes that leverage their adaptive dynamical responses to input electrical signals. They can be viewed as complex physical systems, as they produce higher-order nonlinear dynamics, criticality and phase transitions in the order parameter. These non-trivial dynamics emerge from the interplay between nanoscale transport dynamics (which generate high degrees of freedom in their internal resistance states) and heterogeneous, recurrent network connectivity [24, 31].
Box 1: Memristive devices The ideal memristor is defined by the relationship between flux ($\Phi$) and charge ($q$): $d\Phi=M\,dq,$ (1) where $M$, the memristance, depends on $\Phi$ or $q$. If $M$ is constant, the device behaves as a linear resistor; if $M$ varies, the memristor exhibits nonlinear behaviour. Memristors are typically characterized by a "pinched hysteresis loop" in their current-voltage (II-$V$) graph, which intersects at zero voltage and current [32]. A general description of a memristive device is: $\frac{dx}{dt}=F(x,u),\quad y=H(x,u)u,$ (2) where $x$ represents internal (physical) states, $u$ is the input ($V$ or II), and $y$ the output. The function $H(x,u)$ corresponds to the memristance or memductance [33]. The concept dates back to the 19th century with studies on arc lamps and early radio components. However, in 1971, Leon Chua formalized the memristor as the fourth fundamental circuit element, defining it as a device that “remembers" the history of past inputs. From a practical perspective, it is useful to think of a memristive device as a resistance (e.g. satisfying Ohm’s law $RI=V$) with a time dependent $R(t)$, whose value depends on the history of $V(t)$ (or $I(t)$, depending on the type of device).
This Perspective is intended for researchers across several disciplines. Physicists can view SOMNs as physical substrates to study emergent phenomena [6], like criticality by means of dynamical systems and statistical mechanics frameworks. Material scientists and chemists may be motivated to explore new types of functional nanomaterials with similar properties (e.g. dynamic reconfigurability, long-range spatio–temporal correlations, as well as resistance to perturbations and self–healing, all key properties of adaptive, complex systems [34]). Computational neuroscientists may find value in SOMNs as an abiotic platform for gaining new insights into how coupled, physics–constrained synaptic dynamics produce multi–scale brain-like dynamics and process information [35, 36, 37, 38]. Computer scientists may instead view these physical systems as a potential platform for efficient machine learning applications and for exploring new approaches to solving computing problems.
Beyond scientific interest, SOMNs address practical challenges in embedded edge intelligence [39]. Their ability to learn in situ under resource constraints makes them ideal for edge applications like dynamic sensing, robotics and autonomous systems, as well as continuous real-time monitoring and prediction (see Fig. 1). Indeed, in these edge applications, SOMNs offer a viable solution to AI’s energy crisis [40]. By integrating memory and processing in a single system, SOMNs bypass the von Neumann bottleneck and enable efficient learning directly at the physical interface. To understand how SOMNs can bridge material properties and learning, we examine the physical principles that underpin their brain-like dynamics and highlight how these systems have been implemented to date to perform meaningful learning tasks.
The remainder of this Perspective is organised as follows. sec. 2 discusses the neuromorphic properties and dynamics of SOMNs, including the underlying transport physics that produces physical plasticity (sec. 2.2) and criticality (sec. 2.3). Sec. 3 presents a theoretical explication of SOMNs as complex physical systems, including theoretical methods for elucidating how memristive nonlinearity and memory combined with the physical constraints of Kirchhoff’s circuit laws induce highly non-trivial dynamics in the conductance (sec. 3.1), and how a mean-field model can be constructed (sec. 3.2). Sec. 4 discusses two main approaches to physical learning demonstrated with SOMNs: physical reservoir computing (sec. 4.1) and associative learning (sec. 4.2). We conclude with an outlook in sec. 5.
2 Brain-like dynamics in SOMNs
To understand how SOMNs may be used as physical learning systems, it is first necessary to examine the physical mechanisms behind their nonlinear dynamics.
Nanoscale transport dynamics and a heterogeneous network connectivity underpin the ability of SOMNs to produce emergent brain-like properties, including: plasticity – the ability to adapt to changing input signals; and avalanche criticality – signal propagation exhibiting long-range spatio-temporal correlations [24, 28]. This section describes how these dynamics arise in the context of the most well studied SOMNs, comprised of nanowire (NW) and nanoparticle (NP) networks. Both NW and NP networks show similar brain-like dynamics related to collective memristive switching effects (see Fig. 2). However, the underlying nanoscale transport physics in single NW-NW and NP-NP junctions is notably different.
2.1 Physics of ionic/electronic transport
The dynamics of NW networks are governed by ionic transport in the junctions between NWs. Here, NWs comprised of an electromigrating metal and a dielectric are assembled in a random arrangement where each contact region between two NWs forms a nanoscale electrical junction [41, 42]. Most commonly, NW networks employ silver as the electromigrating metal with either a polymer coating (e.g., Polyvinylpyrrolidone, PVP) [43, 24, 44] and/or metal chalcogenide (e.g., Ag2S [12, 45]), though other materials have also been explored (e.g. Ag-TiO2 [46], AgI [47], Ag2Se [48, 49]). In the nanoscale junctions, the electric field is high enough to drive the dissolution of atomic Ag to Ag+ that then electromigrates towards the opposite wire, where it nucleates back to metallic Ag, forming conductive nanofilaments in the junction [44]. Counter to this additive process, minimisation of the interfacial energy and local electromotive force effects (nanobattery effect) act to dissolve the nanofilaments, becoming the dominant process when the applied voltage is removed. Therefore, similarly to biological neurons where the synaptic strength is associated with changes in ionic flow and to membrane depolarisation and spontaneous repolarisation, NW junctions modulate their conductance depending on input history and are characterised by a short-term memory ranging from milliseconds to seconds. This mechanism is one of the most common memristive behaviours, at the basis of Electrochemical Metallization (ECM) memories and atomic switch devices [50]. In addition to the external electrical bias, the main intrinsic factors are ion mobility and redox rates [51].
Zooming out at the macroscopic level, NW networks are characterised by metastable states induced by the collective interactions among many memristive junctions where nanofilaments are continuously reshaped due to the local redistributions of electrical potential and current. Under voltage sweep stimulation, NW networks exhibit a collective hysteresis response characteristic of memristive systems (cf. Fig. 2(a)).
When large external biases are applied or local inhomogeneities are present, Joule heating can induce the local breakdown of single nanowires, creating gaps where metal ions electromigrate and form nanofilaments, similarly to NW junctions [52]. While the short-term dynamics of such new memristive elements is likely indistinguishable from those of a nanowire junction, the network topology, along with all the local redistributions of electrical potential and current, is irreversibly changed, thus giving to the network a long-term memory characteristic. With a terminology borrowed from the human connectome, the two mechanisms are soundly referred to as reweighting, when an existing synapse (junction) is modulated in conductance, and rewiring, when a new synapse (gap) is created [44].
The conductivity mechanisms in NP networks differ from those in NW networks. There is currently no evidence for electrochemical ionic effects, at least in the most studied system comprised of tin NPs [53, 16] that are typically studied in vacuum. Interestingly, Au discontinuous films  [54] have similar mechanisms, in both air and in vacuum. Electrochemical effects might be expected in, for example, networks of Ag NPs (where electrochemical properties are expected to be similar to that in NW networks) [55], but detailed studies of the mechanisms are still required.
Figure 2: Memristive behaviour of SOMNs. (a) Memristive behaviour of a NW network under voltage sweep stimulation in two-terminal configuration (see inset), characterized by an hysteretic loop in the I-V plane typical of memristive systems. Adapted from ref. [44]. (b) Short-term plasticity effects in a NW network, characterized by potentiation of the conductance state during voltage stimulation followed by spontaneous relaxation after the end of stimulation. Adapted from [56]. (c) Self–similar conductance changes related to memristive switching events in NP networks. Top panel: 2400 s of experimental data; lower panels: successive segments of the data with 10 and 100 greater temporal magnification and greater magnification on the vertical scale. Adapted from Ref. [16].
NP networks are typically fabricated in vacuum thanks to nanoscluster production systems that deposit the nanoparticles on a solid substrate. There they form a 2D layer of connected groups of particles separated by nanogaps[57, 53]. With this method, the surface coverage can be precisely tuned to poise the system near the percolation threshold. This is vital as it results in scale-free network structures and consequently to scale-free dynamics [58]. Interestingly, it has recently been reported [59] that the architectures of 3D percolating networks are substantially the same as those of 2D systems.
Typical metallic materials employed are Sn[57, 53, 60, 16, 61], Bi[62], Au[63, 64, 54, 65], Ag [55, 66, 67, 68, 69, 70], and Cu [71], but recent work in which devices were fabricated using Mo NPs shows the diversity of materials that can be used [72]. The most recent work on Ag devices [69, 70] provides a possible connection to work on NWN devices and suggests that the evolution of Ag network morphology strongly affects memristive / switching behaviour.
The memristive mechanism in NP networks is the following: under an external electrical bias, the strong electric field in the gaps between NPs causes surface atoms to move [73]. At low electric fields this force creates hillocks on the surface of NPs that reduce the gap size, thus modulating the tunneling current [74]. At high electric fields (or if the field is applied for a long time) the hillocks extend all the way across the gaps and form filaments [57]. Once the bias is removed, the hillocks/filaments relax back due to interfacial energy minimisation. The filament formation and breaking processes can be characterised as neuron-like integrate-and-fire mechanisms[75] and timescales are typically in the microseconds range [76]. Filament formation / breaking leads to well-defined changes in conductance i.e. switching events. When one gap switches it causes a redistribution of voltage and current across the network, leading to avalanches of events that resemble those observed in biological neural networks [77, 78] (cf. sec. 2.3).
The proof that the transport is localized into atomic-scale filaments is given by quantum conductance effects, where conductance states are observed in multiples of the fundamental quantum of conductance G0= 2e2/h[57]. Note that quantum conductance effects have also been observed in single NWs [52] and in NW networks [43], as a consequence of the nearly atomic scale of memristive conducting filaments [79]. Note that, as described below, the low field hillock and high field filament modes are useful for different kinds of computation [80, 74, 61, 81, 82, 83, 84].
Thus, charge migration and atomic rearrangement in the nanoscale junctions of NW or NP are responsible for the memristive switching dynamics in SOMNs. These properties result in self-adaptation to changes in the input signals (synapse-like plasticity) and as well as critical behaviour (neuron-like avalanches). These topics are explored in more depth in the next two subsections.
2.2 Plasticity
Plasticity refers to the capacity of a system to modulate its internal states or connectivity in response to external stimuli. In biological neuronal networks, plasticity is fundamental to learning and memory, enabling organisms to retain, discard, and process information in response to new experiences. Such adaptability spans multiple timescales and mechanisms, from transient modulation of synaptic strength (short-term plasticity) to longer-lasting changes that underlie memory formation. Physical systems such as SOMNs can similarly exhibit dynamic reconfiguration in response to stimulation, reflecting not only direct, stimulus-specific conductance changes (analogous to homosynaptic plasticity), but also broader network-level adaptations and structural changes (analogous to heterosynaptic plasticity). These capabilities emerge naturally from the interplay of local transport dynamics and the global architecture of the network.
Following experimental demonstrations of memristive devices as synaptic elements in neuromorphic circuits [85, 86], a wide range of synaptic functionalities has been realised in SOMNs due to their capability to dynamically reconfigure in response to external electrical stimuli, particularly in multiterminal configurations. Experimental results show that the emergent behaviour of NW networks can emulate input-specific changes in synaptic conductance pathways, resembling homosynaptic plasticity. Additionally, multiterminal memristive devices demonstrate that plasticity can also extend to synaptic pathways not directly stimulated, effectively mimicking heterosynaptic plasticity effects [44, 87]. These networks also exhibit short-term plasticity, where synaptic weights—represented by the conductance between neuron-like terminals—depend on recent stimulation history, as evidenced by potentiation–during stimulation followed by spontaneous relaxation after the end of stimulation (see Fig. 2(b)). This effect exploits volatile memristive mechanisms inherent in network elements [44, 88, 31]. Short-term memory effects, experimentally demonstrated to be controllable across multiple timescales [44], enable the emulation of working memory processes observed in biological brains. Working memory includes not only the temporary retention of information but also its active manipulation [88, 89]. Crucially, these short-term memory effects facilitate the physical implementation of reservoir computing paradigms, discussed in sec. 4.1.
Beyond short-term effects, NW networks can also preserve previously established pathways while forming new connections. This capability mirrors metaplasticity in biological systems, where synaptic plasticity depends on the historical context of prior modifications [89]. Furthermore, changes in the NW network’s physical topology, known as structural plasticity, have been observed through electrically induced rupture and electrochemical rewiring of NWs [44, 52]. These long-lasting changes in SOMNs can be exploited to emulate long-term plasticity effects, a crucial aspect for storing knowledge in the system during learning. In this context, using tomographical techniques, it has been experimentally demonstrated that spatially correlated short-term plasticity effects can evolve into long-lasting memory traces resembling biological memory engrams, which represent physicochemical changes underlying stored experiences in the brain [56]. Similarly, short-term and long-term memory functionalities have been reported in NP networks[90]. Collectively, these findings demonstrate the potential of SOMNs as substrates for both processing information and storing knowledge at the material level.
In summary, brain–like plasticity effects exhibited by SOMNs reflects their adaptive dynamics and memory properties. While these properties originate from localised memristive junctions, their dynamic network circuitry results in global reconfigurations and associated memory changes. The following section highlights another example of dynamic reconfiguration exhibited by both biological neuronal networks and SOMNs.
2.3 Criticality
Critical phenomena play a fundamental role in various physical systems, characterised by divergent spatio-temporal correlations and the breakdown of mean-field approximations. Classic equilibrium examples include the paramagnetic–ferromagnetic phase transition described by the Ising model. Importantly, similar critical behaviours emerge in non-equilibrium biological neuronal networks [91], where they are believed to underpin efficient information processing. Here, we highlight that similar criticality has also been experimentally observed in SOMNs, specifically NW and NP networks [16, 75, 92, 93, 94, 95].
NW networks, which are generally fabricated above the percolation density threshold, exhibit criticality when networks are electrically stimulated just above a critical applied voltage [93, 94]. Avalanche criticality observed in Ag-based NW networks is due to electrical percolation when the network dynamics is dominated by electron tunnelling in the junctions [96, 97]. This occurs when junction tunnel gaps form as Ag filaments evolve. These gaps continuously close and re-open as filaments grow and stochastically decay, driving rapid memristive switching between metastable conductance states. At voltages well above the critical voltage, some memristive junctions can form stable filaments that support ballistic transport. When this occurs, NW networks undergo a discontinuous dynamical phase transition resembling dielectric breakdown (also referred to as a “bolt phase”), characterised by the emergence of large amplitude avalanche events [98, 99, 93, 97]. Recent experiments [95] provide deeper insights into critical phenomena in NW networks. These studies demonstrate the programmable induction of critical states via targeted electrical stimulation. Importantly, spatially distinct regions within NW networks can exhibit and sustain local critical dynamics, facilitating spatially correlated interactions that directly influence computational functionality.
While both NW and NP networks show critical dynamics, their measured exponents deviate from known universality classes, hinting at novel behaviour. This may arise due to higher–order interactions or specialised percolation dynamics not fully captured by theoretical percolation models [100, 101], as previously suggested [102, 58, 103]. In particular, the combination of tunneling transport and memristive switching dynamics may distinguish SOMNs from classical percolation universality classes. Recent studies show that tuning network dynamics, via global feedback [104] or local control of junctions [95], can steer systems toward different universality classes, offering pathways to control and apply criticality in SOMNs.
Avalanche criticality in SOMNs is an important concept, as it is a prominent feature that mirrors experiments performed in neuroscience [105]. It is in fact hypothesised that functioning near this point optimises information processing [77, 78, 91]. SOMNs are controllable and experimentally accessible physical systems, and offer valuable abiotic platforms for rigorously testing this hypothesis through direct implementation and analysis of computational tasks (discussed further in sec. 4).
3 SOMNs as dynamical complex systems
As highlighted in the previous section, SOMNs are complex and dynamic electrical circuits that produce emergent nonlinear dynamics. Here, we present a theoretical framework that explicitly treats SOMNs as dynamical complex systems subject to the physical constraints of Kirchhoff’s electrical conservation laws. This theoretical framework enables models to be developed to gain deeper insights into how the experimentally observable collective dynamics arise from the memristive network interactions.
In particular, these methods are useful for identifying dynamical transitions and to characterise the dynamical regimes that may enhance learning [106, 107, 108], as elaborated in sec. 4.
3.1 Theoretical Methods
A practical theoretical framework for modelling SOMNs is the lumped circuit approximation. By representing their conductance dynamics using memristive elements, with voltage drops across the junctions, SOMNs can be modelled effectively as assemblies of interacting subcircuits, each corresponding to an individual memristive junction (neglecting the resistance of individual NWs or NPs). Their interactions are governed by Kirchhoff’s laws, resulting in complex, high-dimensional nonlinear dynamical systems.
A convenient way to mathematically capture the diverse nonlinear dynamics observed experimentally is to model each junction’s conductance as a dynamical variable dependent on its voltage or current history. This approach provides a simplified, circuit-inspired model that captures memory and adaptation using only voltage, current, and time, e.g. quantities that are physical observables. This method applies to both NW [109] and NP [110] networks and can be extended to other systems [111].
The internal state of each memristive junction typically evolves according to a first-order differential equation of the general form (cf. Box 1):
$\frac{dg}{dt}=F\bigl{(}g,\,\delta v(t),\,t\bigr{)},\quad G=G(g),$ (3)
where $g\in[0,1]$ represents a normalized internal memory parameter, and $v(t)$ denotes the local instantaneous voltage across the junction. The function $F$ encodes how the conductance changes in response not only to $v(t)$ but also to make explicit time dependence, allowing modelling of more complex memory behaviours. Equation $G(g)$ represents the electrical conductance of the junction in terms of $g$. An equivalent representation can also be made in terms of electrical resistance.
While the lumped-element model introduced above is widely applicable (particularly for NW networks), alternative modeling approaches have been developed for NP networks, where different memristive mechanisms are at play. For instance, a potentiation–depression rate balance equation, originally derived from detailed studies of single memristive junctions [112], has been used to describe dynamic conductance changes at both the junction level [27] and in spatially extended networks via a grid-graph modeling framework [87]. These rate-based models offer a complementary theoretical perspective by naturally capturing key plasticity phenomena, including short-term plasticity, metaplasticity, and structural plasticity, as emergent from cumulative electrical activity. As such, they are especially well-suited for describing the temporal evolution of NP networks, where conductance modulation often involves field-induced atomic rearrangement rather than ionic transport mechanisms typical of NW systems.
The lumped-element approximation simplifies the system by reducing each junction to a single memory variable, omitting other physical degrees of freedom that may affect network dynamics. Small but finite capacitive effects at the junctions, though often excluded from basic memristive models, have been experimentally observed and may become significant under certain conditions [113, 114]. These effects introduce distinct memory-like dynamics and can be examined using techniques such as Nyquist impedance spectroscopy.
From a practical standpoint, Kirchhoff’s laws can be numerically solved for a given network and initial junction states to determine voltages and currents across each element. Integrating the dynamical equations (3) then reveals the time evolution of the conductances. This direct approach is widely used in the literature [75, 110, 27, 93, 74] to assess the effective conductance and functional behaviour of SOMNs.
To expose how topology and constraints shape dynamics and to gain analytical insight, an analytical method using the projector operator formalism (see  Box 2) has been developed[115]. This incorporates Kirchhoff’s constraints into the junction evolution equations via projection matrices: $\Omega_{B}$ in the nodal formalism and $\Omega_{A}$ in the mesh formalism. Within this framework, the network dynamics reduce to a system of coupled nonlinear ordinary differential equations:
$\frac{d\vec{g}}{dt}=\vec{\mathcal{F}}\bigl{(}\vec{g},\,\vec{\delta v(t)},\,\Omega_{B},\,t\bigr{)},$ (4)
where $\Omega_{B}$ ensures explicit enforcement of Kirchhoff’s laws.
In equations like (4), nonlinear memory effects emerge from the interplay between network topology and time-dependent junction conductances. Unlike resistive networks with fixed internal states, memristive networks adapt their conductance (or resistance) in response to external inputs. The topology is captured via projector matrices, $\Omega$, which typically appear in the equations when Kirchhoff’s laws are implemented analytically also via matrix inverses of the form $(I\pm\chi\Omega\mathcal{G})^{-1}$, where $\mathcal{G}_{ii}=g_{i}$ is a diagonal matrix [115, 109].
This formalism can be extended to more general network motifs, where edges represent subcircuits with complex behaviour, such as memristive, inductive, or capacitive coupling [116]. Though not explored in detail here, such extensions may be important for accurately modelling phenomena like the experimentally observed capacitive effects in SOMNs [113, 114]. Crucially, the matrix inverse structure in the projector formalism induces nontrivial feedback loops within the network. These loops underpin emergent behaviours such as self-organisation and dynamical phase transitions [109], providing a unified analytical framework for linking local memristive dynamics to rich global behaviour.
Incorporating stochasticity into SOMN dynamics offers a powerful extension of the theoretical framework. Stochasticity is a fundamental characteristic of NP networks [75], particularly in the high-voltage regime where the formation and rupture of conductive filaments leads to continuous redistributions of voltage and current, thereby triggering new switching events. The inherent complexity of these networks [58] further implies that the timing of such events is effectively random. Moreover, the filament formation process itself is intrinsically stochastic [76]. Accordingly, SOMNs can be modelled as stochastic dynamical systems [30], with network conductance described, for example, by an Ornstein–Uhlenbeck process that blends stimulus-driven evolution with intrinsic fluctuations. This stochastic mean-field approach captures features such as conductance noise and jumps, while enabling analysis of the potential landscape and the interplay between noise, steady states, and computational capability. In this way, stochastic models provide deeper insight into collective behaviour that cannot be fully described by deterministic frameworks.
The “solid–liquid brain” metaphor [117] illustrates the adaptability of SOMNs: unlike static resistive networks (“solid brains”), SOMNs act as “liquid brains”, flexibly adapting to stimuli, like biological systems [118, 119]. Moreover, equations of the form (4) incorporate memory and topology while enabling Lyapunov function construction, allowing stability analysis akin to Hopfield networks. These functions reveal attractor dynamics in a potential landscape [120], underpinning computational features such as nonlinear responses, metastability, and configuration exploration, which are key for optimisation tasks.
3.2 Mean field theory of conductance transitions
Critical phenomena in SOMNs are closely linked to self-organisation and feedback dynamics, which enable adaptive states under external driving. These behaviours often appear as scale–free event distributions—such as avalanches [16]—and are key to understanding emergent correlations. In SOMNs, global conductance can exhibit sharp transitions with voltage or component density, arising from the interplay between local dynamics and global connectivity. The mechanisms behind these critical states warrant further theoretical investigation. Related studies in neuromorphic nanowire networks [93, 94, 95], percolating NP networks [75, 54], and stochastic switching systems [76] provide experimental and computational evidence for avalanche dynamics and phase transitions. Theoretical insights into such transitions have also been drawn from analogies to random fuse models [121], dielectric breakdown [98], and tunneling-dominated percolation [102], which are relevant to understanding the observed deviations from classical universality classes in SOMNs [58, 101].
Mean-field models have been developed for the percolation transition [99], drawing on approaches from random fuse models [121]. While such mean-field approaches offer insight into general system behaviour, they fall short in accurately predicting critical exponents, which require more refined analysis [122] and consideration of effects like tunneling that are normally not consisdered in percolation models [102]. Crucially, to be effective, mean-field theory must incorporate the “liquid” nature of SOMNs—their dynamic, state-dependent connectivity, which is central to capturing their collective and emergent properties.
Figure 3: Modeling emergent behaviour of memristive networks. (a) Experimental data on potentiation and relaxation dynamics related to short-term plasticity effects in NW networks under voltage stimulation (blue line), grid-graph modeling (black dashed line) and mean-field theory results (green dashed line) obtained by using methods outlined in sec.  3. (b) Results from grid-graph modeling allow to visualize spatiotemporal dynamics of the network, characterized by the progressive formation of a conductive pathway bridging stimulated electrodes that progressively fades way after the end of stimulation. The red intensity is proportional to the conductance of edges, while the blue intensity is proportional to the node voltage. Adapted from Ref.  [123] and  [109].
[…]
Using the theoretical framework outlined in sec. 3.1, recent work has formulated a dynamical mean-field theory that attempts to characterize the voltage-induced dynamics of SOMNs by reducing equations of the type (4) to a single mean field dynamical equation. This is written in terms of an effective potential $\mathcal{V}_{\Delta v}(\langle g\rangle)$, which governs the evolution of the network’s average conductance state [120, 109], and allows determination of whether the interacting system is in the low or high conducting phase. These phases are determined by the state of the junctions, e.g. all on or all off. The resulting equation of motion is of the form:
$\frac{d\langle g\rangle}{dt}=-\frac{d}{d\langle g\rangle}\mathcal{V}_{\Delta v}(\langle g\rangle),$ (7)
where $\Delta v$ is the applied voltage. As voltage increases, the shape of $\mathcal{V}$ changes, potentially introducing multiple minima corresponding to different conducting states. A mean field equation of the type (7) has been validated experimentally [109] and Fig. 3 shows a comparison between experimental data on short-term synaptic plasticity effects and theoretical response curves obtained through mean-field theory, as well as simulation results based on a grid–graph model that allows to directly visualize spatiotemporal dynamics of the system [123].
Figure 4: Conductance transitions in memristive networks. a. Schematic illustration of conductance phase transitions in SOMNs, where the macroscopic order parameter (average conductance $G$) depends on control parameters such as applied voltage $V$, nanowire/nanoparticle (areal) density $\rho$, or average connectivity $d$. The shaded region represents the transition regime where emergent collective behaviour manifests. Depending on system parameters and underlying dynamics, this transition can be continuous (second-order), abrupt (first-order), or gradual, reflecting different types of dynamical phase behaviour. b. Numerical simulation showing the time evolution of internal junction conductance states (red lines) across the network following a voltage step near the critical point, when the effective potential develops multiple competing minima. The black line shows the ensemble average $\langle g\rangle$ across all junctions, while the dashed blue line corresponds to the prediction from mean-field theory (MFT). The divergence between simulation and MFT near the critical regime illustrates the breakdown of mean-field approximations due to the onset of strong feedback and topological reconfiguration. The inset shows the subcritical regime (single minimum in the effective potential), where junction conductances decay smoothly (laminar dynamics) and MFT matches simulations well. These results highlight the emergence of collective, nonlinear dynamics and the importance of including feedback and stochastic effects beyond mean-field approaches.
Various methods have been used to obtain an equation of the form (7) from eqn. (4). At the core of these methods is how the matrix inverse $(I\pm\chi\Omega_{B}\mathcal{G})^{-1}$ enters into eqn. (4). This matrix inverse, once expanded, leads to non-linear terms between the variables $g_{i}$’s, and thus couples the junctions memory states across the network, and its structure gives rise to feedback and phase transitions. The first method is based on random matrix theory averaging of the inverse while keeping fixed row and column sums  [120, 128, 129]. Another method involves preserving certain graph-theoretic properties (e.g., number of cycles), leading to orthogonal transformations $\Omega_{B}^{\prime}=O\Omega_{B}O^{T}$ (with $O$ an orthogonal matrix), enabling exact averaging over transformations via Weingarten calculus [130, 131]. Alternatively, and more brutally in [109], the mean field inverse has been obtained by minimising the Frobenius norm between the exact and the mean field inverse. The resulting effective potential shape depends on the voltage, which implies that conductance transitions between the low and high conductances are controlled by a macroscopic theory. When there is only a single minimum, the dynamics of the junction is laminar and the MFT works well numerically. However, the devil is in the details. Near transition points, e.g. when the effective potential develops multiple minima, the system becomes susceptible to noise, which can be amplified and can induce transitions between these two states (cf. Fig. 4b), and a failure of the MFT. Depending on the shape of the effective potential, these transitions can be both of the first or second order (continuous or discontinuous) type, dependent on voltage, in addition to static geometrical effects due to density (cf. Fig. 4a). Further studies and new techniques [132] are necessary to understand the behaviour of SOMNs as complex physical systems. It is important to stress that mean field theory is not effective near dynamical phase transitions; as we can see in Fig. 4b, the match between the dashed blue (MFT) and solid black line (simulations) is poor near the threshold between the laminar and turbulent regime, while it is quite remarkable in the laminar one (inset, subcritical). Thus, alternative approaches (e.g. self-organised criticality [133]) are necessary to describe such critical points.
In summary, theoretical models provide a foundation for understanding how SOMNs generate the collective nonlinear dynamics observed experimentally. From lumped-element to mean-field approaches, these frameworks reveal how memory, phase transitions, and adaptive responses emerge. These emergent dynamics represent physical information encoding and form the basis of a physical learning substrate, especially near transitions. The next section explicitly shows how these dynamics enable learning.
4 Physical learning with SOMNs
Figure 5: Physical learning paradigms for self-organising memristive networks. Arbitrary input data (left) is first pre-processed and encoded into input voltage signals (centre left) and delivered to a multi-terminal SOMN device (centre main), where conductance changes adapt to the input signal patterns. Dynamical reconfiguration of the SOMN, governed by internal transport dynamics and Kirchhoff constraints, physically compresses signal information to enable learning (output). Two representative learning paradigms are shown (right). In reservoir computing, the SOMN’s high-dimensional dynamical features are read out (red solid curves) and exploited in an external machine learning model; feedback to the input (red dashed curve) is also applied in some cases (e.g. recursive prediction). In associative and contrastive learning, feedback is applied based on output (green dotted lines). Both approaches enable flexible, substrate-intrinsic learning using only input-output access, without any direct addressability of memristive components.
With an understanding of the mechanisms that underlie adaptive dynamics in SOMNs, we now show how these properties can be operationalised in physical learning systems. To provide some context, it is important to note that solving complex learning tasks requires nonlinear integration, as became apparent after the perceptron was introduced  [9, 134]. Nonlinear integration can be achieved either through multilayer networks [11] and/or recurrent dynamics [10].
Neuromorphic approaches aim to realise such nonlinearities by mimicking biological synapses and/or neurons, either through emulation or simulation [17]. The emulation approach involves exploiting material and devices (e.g. memristors) that natively exhibit neuromorphic nonlinear dynamics [135]. The simulation approach instead typically involves algorithms that simulate point neuron spike models [136]. While both approaches can mimic synaptic plasticity, which is generally regarded as the key mechanism underlying the brain’s unique learning abilities [137], how this is implemented at the circuit device level differs considerably. For example, synaptic conductances in memristor crossbar devices are often used to physically instantiate matrix–vector operations to perform inference on globally-trained ANN models. On the other hand, spike-based neuromorphic circuit devices typically implement local learning algorithms (e.g. spike-timing dependent plasticity, STDP).
Here, we highlight how physical learning can be achieved with SOMNs by exploiting their local and global brain-like nonlinear dynamics (including memory), without requiring direct addressability or control of their memristive circuit components. Two key learning paradigms are described below and illustrated in Fig. 5: physical reservoir computing and associative learning. These learning paradigms demonstrate how the dynamical states of SOMNs, as complex physical systems, can be exploited for learning either in an external output layer, in the case of reservoir computing, or within the network itself, in the case of associative learning.
4.1 Physical reservoir computing
Reservoir computing is a machine learning paradigm where a reservoir (a recurrent neural network) maps an input signal into a higher dimensional manifold that is then analysed by an output layer, which is the only part that needs to be trained [138], making this an efficient learning scheme (the weights inside the neural network are fixed and not trained.)
In the physical or in materia implementation, a non-linear physical system is exploited as the reservoir [139, 140, 141, 142, 143]. In this context, SOMNs have been demonstrated to satisfy the key requirements of a physical reservoir [47, 106, 45, 144, 145, 31, 123, 146, 80, 74, 81, 147], endowing i) high dimensionality, with internal dynamics made accessible via multiterminal device configurations; ii) high–order nonlinearity, realised through atomic / ionic transport mechanisms and recurrent network structure; iii) fading memory property (to ensure that reservoir states rely only on recent–past inputs and are independent of distant–past inputs) [106], as discussed in Sec. 2.2.
To date, physical reservoir computing with SOMNs has been experimentally demonstrated in a wide range of computing tasks. One of the earliest demonstrations was waveform regression using NW networks [45]. For NW-based SOMNs, these also include pattern recognition [123, 48], audio classification [47, 49], sequence memory [48] and waveform regression [45]. For SOMNs based on NPs physical reservoir computing has demonstrated time series prediction, non-linear transformation, waveform discrimination, memory capacity and speech recognition [74, 146], as well as speaker and Braille recognition [147]. Time series prediction and speech recognition was also experimentally achieved on NW-based SOMN in [148, 149], and theoretically investigated in [106]. In addition, the experimental results have been complemented by simulations that have been used to test in silico a range of issues including the scalability of these systems, the architecture of the reservoir[150, 74, 103], the impact of the positioning [92] and the number of electrodes and the dimensionality of the reservoir[110]. A wide range of computing tasks have been simulated including pattern recognition  [123, 151, 144], time series prediction  [149, 123, 145, 152], speech recognition  [148], regression [93, 31, 153] and to learn chaotic dynamics  [154]. Furthermore two-dimensional tracking and autonomous swarm generation has also been simulated [81].
The internal states of SOMNs and their emergent spatiotemporal dynamics [116] are both measured via multiterminal electrical measurements  [155, 61, 156]. These measurements are typically performed by measuring output currents or floating voltages, or by constructing a conductance matrix where each element corresponds to the conductance between a pair of electrodes. Reservoir computing can also be implemented in a two-terminal memristive network by adopting a time multiplexing scheme, as demonstrated for both NP [74] and NW networks [30].
Notwithstanding these attributes, physical reservoir computing is significantly different from algorithmic reservoir computing in the following ways: (i) physical reservoirs are constrained by physics and in the case of SOMNs, Kirchhoff’s electrical circuit conservation laws; (ii) SOMNs produce collective memory dynamics by virtue of their memristive elements; and (iii) physical reservoirs are dynamic networks, with non–static connectivity [157] For these reasons, physical reservoir computers cannot be mathematically described in the same way as algorithmic reservoir computers, which apply a mathematical nonlinear activation function solely to nodes in a recurrent network with a random, static connectivity structure, with many hyperparameters requiring optimisation. However, dynamics typical of SOMNs can be leveraged to develop new neural network models based on the reservoir computing paradigm [158, 154].
Similar to biological systems where functionalities are inherently related to the topology of neuronal networks [159], higher reservoir computing performance can be achieved using SOMNs with scale-free topology [74]. Higher computational performance can also be achieved experimentally via parallel reservoir architectures, which increase the reservoir dimensionality and richness of its output dynamics [146]. Importantly, it has been shown through simulations that performance in RC tasks using NW networks contacted with a restricted number of electrodes is close to the upper bounds achievable when using information from every NW [110], a key consideration for the experimental realisation of these systems, where information from each NW is not practically accessible. Such simulations provide additional insight into experimental findings that a large number of electrode readouts may not be necessary to achieve high performance, at least on some RC tasks implemented on NW network devices [155, 48]. Simulations and experimental implementations of more complex RC tasks are needed to qualify these results further.
Additionally, it has been shown experimentally that the spatiotemporal information processing capabilities in multiterminal systems allow unconventional implementation of RC by exploiting the same electrodes as both reservoir inputs and readouts, allowing a reduction of hardware complexity without limiting computing capabilities [151]. Additionally, while the output layer is typically trained in software,
a fully memristive architecture can be realised by coupling SOMNs acting as a reservoir with memristive crossbar arrays [123, 160]. Furthermore, it has been revealed that memory patterns embedded in dynamical features of NW networks enable online learning and recall of spatiotemporal sequence patterns, thus showing the possibility of online learning from rich spatiotemporal dynamics generated by the SOMN, in contrast to data-driven statistical machine learning techniques [48].
Several open questions remain in physical reservoir computing with SOMNs, including the final energy consumption for real world tasks, and whether methods can be developed for more self-contained learning that does not rely on training an external layer [2]. Indeed, such an approach could help bridge the gap between physical reservoir computing and associative learning, which is discussed in the next section.
4.2 Associative learning
Associative learning, i.e. the formation of associations between stimuli and responses, can be implemented in SOMNs by taking advantage of their adaptive conductance changes (i.e. plasticity, cf. sec. 2.2). External electrical stimuli applied to SOMNs drive local conductance changes that can lead to global reconfigurations in network states as unique conductance pathways are established for different patterns in the stimuli [161, 46]. The memory encoded in their conductance states [56] enables SOMNs to recall, or self–learn, these input–output associations. In contrast to reservoir computing, learning is intrinsic to the network: repeated stimulation leads to material-level changes in conductance pathways, enabling SOMNs to self-organise and recall associations without external training mechanisms. This is more analogous to biological synaptic plasticity than statistical optimisation. Teaching SOMNs input–output relationships via feedback, without requiring explicit programming also contrasts sharply with conventional implementations of physical neural networks, which typically rely on direct internal state addressability, for example, to physically implement Hopfield associative memory networks [162].
As self–adjusting to external stimuli mimics biological learning [35, 37], associative learning in SOMNs effectively provides a physical and material basis for cognition [28, 163]. Experimental and theoretical studies have demonstrated the capacity of SOMNs to emulate core cognitive functions, including associative memory and learning [163, 161] as well as working memory [89]. Figure 5 schematically illustrates the learning loop: input stimuli perturb the SOMN; readouts are compared to targets; and corrective inputs guide reconfiguration through internal plasticity [108]. Moreover, implementation of the nn–back working memory task has revealed that short–term memory can be consolidated into long–term memory by applying both positive and negative feedback, in a manner similar to contrastive learning [89, 35, 164]. This metaplasticity is enabled by the heterogeneous and recurrent network connectivity [56].
Both positive and negative feedback enables SOMNs to adapt their electrical connectivity and conductance paths iteratively, to promote the emergence of robust functional mappings. This iterative correction process can be viewed as a form of evolutionary search, in which input–output feedback functions as a fitness signal, guiding the system toward functional configurations [165, 166]. In this view, learning resembles an adaptive exploration of the network’s dynamical state space, governed by error correction, constrained by physical laws and driven by environmental cues [37]. These mechanisms underscore that learning is not defined by task complexity, but rather by the system dynamics and efficiency of feedback. SOMNs bridge conventional and biological models, offering real–time adaptability, dynamic reconfiguration, and connectivity–driven behaviour [35, 37]. A key open question is how many states can be learned and/or stored in these types of physical architectures?
5 Outlook
The brain demonstrates that intelligence can emerge from the physical properties of matter. This realisation continues to motivate the search for materials and architectures with innate brain-like physical properties that may thus be considered promising substrates for emulating core cognitive functions such as memory and learning. Self-organising memristive networks are compelling candidates in this endeavour. As highlighted in this Perspective, the emergent nonlinear dynamics produced by these non-equilibrium physical systems can be exploited in adaptive learning paradigms, without the need for explicit programming. This sets the stage for developing broader technological applications and exploring long-term implications.
Central to their role in physical learning systems is the emergence of global, collective dynamics from relatively simple, local interactions. Nanoscale transport processes (i.e. electron and ionic migration, and atomic rearrangement) continuously modulate junction-level conductance. These local dynamics propagate globally under Kirchhoff constraints and the complex network topology gives rise to dynamical phase transitions, a hallmark of emergent phenomena. This interplay between local plasticity and global self-organisation underpins the functional properties that serve to encode information and enable learning.
Here, we have highlighted two learning paradigms, reservoir computing and associative learning, that leverage the internal physics of self–organising memristive networks. A range of other schemes are now under development, such as those based on spiking computation[82], probabilistic computing [83, 84] and energy–based algorithms (e.g. Hopfield networks, Boltzmann machines) that use binary units which may be physically instantiated with thresholded memristive switching. Recently it was shown that SOMNs that incorporate both spiking neural characteristics and memristive synapses allow new types of learning behaviour [167]. Another research question that arises is: what determines the upper bound of the ability of these memristive networks to process information (i.e. compression, coding/decoding of data)? In some cases, it may be possible to minimise the amount of external computation required to obtain a desired output, effectively making the learning training–free and independent of learning algorithms [168].
From a theoretical standpoint, progress in modeling memristive networks, using tools such as graph-theoretic projector operators or mean-field approximations, continues to enhance our understanding of their functional capacity. These models enable us to describe and even guide the evolution of useful dynamical states, allowing for a form of guided self-organisation [169]. By tuning network density, topology, or stimulation protocols, one can steer the system toward regimes that balance flexibility, stability, and computational richness. On a pragmatic level, recent modelling work has demonstrated the way that different memristor types affect reservoir computing performance [170]
Altogether, self-organising memristive networks represent a material route toward physically embodied intelligence. They shift the paradigm from computing architectures that are engineered and explicitly programmed, to systems that are grown and trained through physical interaction and feedback. By embedding function into structure, and adaptation into dynamics, these materials may offer a path toward scalable, autonomous learning platforms that mirror both the form and function of biological cognition. Taken together, this convergence of physical dynamics and computational function positions SOMNs not merely as neuromorphic analogs, but as a distinct, material foundation for intelligent physical systems.
Ultimately, such systems suggest a future where intelligence is not imposed onto matter, but emerges from it; where computation is not merely symbolic, but physical; not strictly deterministic, but adaptive; and not rigidly encoded, but continuously evolving. In this framework, self–organising memristive networks are a resource that can be harnessed for a new generation of brain–inspired physical learning architectures.
Acknowledgements. F.C.’s work was conducted under the auspices of the National Nuclear Security Administration of the United States Department of Energy at Los Alamos National Laboratory (LANL) under Contract No. DE-AC52-06NA25396, and supported financially via DOE LDRD grant 20240245ER. S.A.B acknowledges support from the MacDiarmid Institute for Advanced Materials and Nanotechnology and the Marsden Fund NZ. G.M and C.R. acknowledge funding by NEURONE, a project funded by the European Union - Next Generation EU, M4C1 CUP I53D23003600006, under program PRIN 2022 (prj code 20229JRTZA). G.M. acknowledges funding by the European Union (ERC, "MEMBRAIN", No. 101160604). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.